[Brahms-l] Shift report 20040111 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 11 2004 - 16:00:33 EST
------------------------------------------------------------
Date: Sun Jan 11 16:00:33 2004, Shift: 08:00-16:00
Supervisor: Oana Ristea, with: Ionut Arsene
Log book#15 pages 30 - 30

Shift Summary:
------------------------------------------------------------
1. Beam from the beggining of the shift until 9:00. After that there was three unsuccesfully attempts to get beam.

2. 9550 - Data run with FS@4deg (1/sqrt{2}A) and MRS@90 deg (1050A) - 126k events for trig.6
   9551 - Non-TPC pedestal run
   9552 - TPC pedestal run

3. H2 TDC Top25 not working.
    
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 11 16:01:27 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 11 2004 - 16:01:36 EST